# Structured Learning of Compositional Sequential Interventions

### v1

@@ Appendix D More Results
In this section, we present more experimental results on our method. This includes: (1) the effect of choosing $r$; (2) further examples of visualization and empirical examples.
D.1 Effect on Choice of r
The effect of different $r$ is presented in Fig. 3, with $r=5$ being the ground truth parameter which data is generated. We have also plotted the performance of our method under $r=3$ and $r=10$. We observed that when $r$ is smaller than the ground truth, the model tend to under-fit the data with a slight higher MSE error. When $r$ is bigger than the actual value, the MSE error can be further reduced but also at the same time results in very large variance. We also show that this can be mitigated with further regularization such as applying $L_{1}$ and $L_{2}$ terms during likelihood learning process.
The insight we drawn from this is that when apply our approach in real-world problems, we could potentially choose a higher $r$ without much concern about what an exact value should be. Using a large $r$ leads to higher variance but in general performs reasonable well compare with knowing the true $r$ (recall that here $r=5$). We also notice that applying regularization technique even when we know the right dimension of $r$ can be beneficial, leading to more stable performance (with lower variance), but does not change much of the mean estimate as shown in Fig. 3. This maybe partially caused by the gradient optimization process when training the deep neural network. Hence, we claim that the main price to be paid for a large choice of $r$ is that this makes data requirements stricter (such as the number of steps prior to the first intervention) in order to guarantee identification.
@@ D.2 Additional Figures
Figure 5: Model-based uncertainty quantification for Reconstruction (CSI-VAE-1).
Figure 6: Demonstration of prediction for synthetic data for CSI-VAE-1.
Figure 7: Residual distribution of our problems (based on CSI-VAE-1 model). In general, we observe a very long tail effect across model predictions.
@@ Appendix E Uncertainty Quantification
For the sake of illustration, we generate another new dataset with a random seed of $42$ using the synthetic simulator, based on the following specifications: $r=5$, $T=20$, $T_{0}=10$, $K=5$, $N=50000$, $z_{\text{dim}}=5$, $I=3$. We also generated two further datasets of $M=2000$, one for calibration and one for test.
@@ E.1 Plug-in Model-based Uncertainty Quantification
We first present the model based uncertainty quantification at level of $\alpha=0.95$, in Fig. 8. Given our conditional mean and homoscedastic error variance $\sigma^{2}$ estimates from data, we can calculate predictive intervals as $(\mu-1.96\times\sigma,\mu+1.96\times\sigma)$. For purely model based uncertainty quantification with plug-in parameter estimates, we do not need an additional calibration dataset, but it is an obviously problematic one as it does not take into account estimation error and sampling variability. The coverage rate is $57.75\%$ on the test dataset without further calibration.
@@ E.2 Conformal Prediction
We then present the model-free uncertainty quantification at level of $\alpha=0.95$ (based on the Setup 1 in section 3.2), shown in Fig. 9. We use the absolute value over the predictive residual as our conformity score function, $\text{abs}(y-f(x))$. We calibrate our model output based on the calibration dataset and then apply it on the test dataset. We obtained a coverage rate of $95.30\%$. This is a significant improvement upon what we see in Fig. 8 and demonstrates the effectiveness of the conformal prediction approach.

### v2

@@ Appendix D More Results
In this Section, we present more experimental results for our method. This includes: (1) the effect of choosing $r$; (2) further examples of visualization and empirical results.
D.1 Effect of the Choice of $r$
The effect of different choices of $r$ is presented in Fig. 3, with $r=5$ being the ground truth parameter from which the data is generated. We have also plotted the performance of our method under $r=3$ and $r=10$. We observed that when $r$ is smaller than the ground truth, the model tend to under-fit the data with a slightly higher MSE error. When $r$ is bigger than the actual value, the MSE error can be further reduced, but resulting in very large variance. We also show that this can be mitigated with further regularization, such as applying $L_{1}$ and $L_{2}$ penalty terms during the likelihood learning process.
The insight we draw from this is that, when applying our approach to real-world problems, we could potentially choose a higher $r$ without much concern about what an exact value should be. Using a large $r$ leads to higher variance but in general performs reasonably well compare to knowing the true $r$ (here, $r=5$). We also notice that applying regularization techniques, even when we know the right dimension of $r$, can be beneficial. This leads to a more stable performance (that is, with lower variance), but does not change much of the mean estimate as shown in Fig. 3. This maybe partially caused by the gradient optimization process when training the deep neural network. Hence, we claim that the main price to be paid for a large choice of $r$ is that this makes data requirements stricter (such as the number of steps prior to the first intervention) in order to guarantee identification.
@@ D.2 Additional Figures
Figure 5: Model-based uncertainty quantification of reconstruction for CSI-VAE-1.
Figure 6: Demonstration of prediction, in the synthetic data case, for CSI-VAE-1.
Figure 7: Residual distribution examples for CSI-VAE-1. In general, we observe a very long tail effect across model predictions.
@@ Appendix E Uncertainty Quantification
For the sake of illustration, we generate another new dataset with a random seed of $42$ using the synthetic simulator, based on the following specifications: $r=5$, $T=20$, $T_{0}=10$, $K=5$, $N=50,000$, $z_{\text{dim}}=5$, $I=3$. We also generated two further datasets of $M=2,000$, one for calibration and one for testing.
@@ E.1 Plug-in Model-based Uncertainty Quantification
We first present the model-based uncertainty quantification at level of $\alpha=0.95$, in Fig. 8. Given estimates of the conditional mean and homoscedastic error variance $\sigma^{2}$ learned from data, we can calculate predictive intervals as $(\mu-1.96\times\sigma,\mu+1.96\times\sigma)$. For purely model based uncertainty quantification with plug-in parameter estimates, we do not need an additional calibration dataset, but it is an obviously problematic one as it does not take into account estimation error and sampling variability. The coverage rate is $57.75\%$ on the test dataset without further calibration.
@@ E.2 Conformal Prediction
We then present the model-free uncertainty quantification at level of $\alpha=0.95$ (based on the Setup 1 in Section 3.2), shown in Fig. 9. We use the absolute value of the predictive residual as our conformity score function, $|y-f(x)|$. We calibrate our model output based on the calibration dataset and then apply it on the test dataset. We obtained a coverage rate of $95.30\%$. This is a significant improvement upon what we see in Fig. 8 and demonstrates the effectiveness of the conformal prediction approach.